# Agent Design Pattern Catalogue: A Collection of Architectural Patterns for Foundation Model based Agents

### v1

@@ A Collection of Architectural Patterns
Foundation model-enabled generative artificial intelligence facilitates the development and implementation of agents, which can leverage distinguished reasoning and language processing capabilities to takes a proactive, autonomous role to pursue users’ goals. Nevertheless, there is a lack of systematic knowledge to guide practitioners in designing the agents considering challenges of goal-seeking (including generating instrumental goals and plans), such as hallucinations inherent in foundation models, explainability of reasoning process, complex accountability, etc. To address this issue, we have performed a systematic literature review to understand the state-of-the-art foundation model-based agents and the broader ecosystem. In this paper, we present a pattern catalogue consisting of 16 architectural patterns with analyses of the context, forces, and trade-offs as the outcomes from the previous literature review. The proposed catalogue can provide holistic guidance for the effective use of patterns, and support the architecture design of foundation model-based agents by facilitating goal-seeking and plan generation.
@@ 1 Introduction
  • Reasoning uncertainties embedded in the underlying foundation models due to hallucinations. It is difficult for foundation models to fully understand and perform complex tasks, thence agents may generate inaccurate responses. The challenge can result in the issue of probability of success across a long chain of steps in a plan. Each step may have a reasonable success rate, while slightly adjusting a few steps can dramatically affect the overall success rate.
  • Limited explainabilty as the sophisticated internal architecture of foundation models make them “black box” to stakeholders. Hence, agents often struggle to interpret their reasoning steps, which may affect the reliability, robustness and the overall trustworthiness of agent systems.
  • Complicated accountability process considering the interactions between different stakeholders, FM-based agents, non-agent AI models, and non-AI software applications in the overall ecosystem. Users ask an agent to achieve certain goals, while the agent can assign tasks to other agents or leverage external tools. The complexity stems from the fact that accountability usually cannot be delegated while responsibility can be shared from one entity to another.
In this regard, we present a catalogue of patterns for the design of foundation model-based agents in this paper. Please note that “agent” can refer to i) AI acting on behalf of a person based on strictly predefined rules with no goal-seeking or planning involved, or; ii) AI that can devise the sub-goals from a high-level goal provided by users. Hereby, we claim that in this study, we focus on the second concept of “agents” that are capable of goal-seeking and plan generation. In software engineering, an architectural pattern is a reusable solution to a problem that occurs commonly within a given context in software design. Our pattern catalogue includes 16 patterns that were identified based on the study conducted by Lu et al. [3]. The intended audience of collected patterns is software architects and developers who are interested in FM-based agent design and implementation. The contributions of this paper include:
@@ 2 Background & Related Work
The introduction of OpenAI’s ChatGPT [4] in November 2022 has gathered over 100 million users in two months upon its release, becoming the fastest-growing consumer internet app of all time [5]. This has also initiated the race among big tech arms in the development of FM and GenAI products. For instance, Google rolled out its own GenAI product, the Bard models  [^3] https://ai.google/static/documents/google-about-bard.pdf then released Gemini  [^4] https://gemini.google.com/. Anthropic has also emerged as one of the major LLM providers since the launch of their Claude models  [^5] https://www.anthropic.com/claude There are also many open-sourced FMs, such as Llama  [^6] https://llama.meta.com/ and Mistral  [^7] https://mistral.ai/.
With the explosive growth of FMs, it is highly notable that FM-based agents come into the picture. AI agents are typically designed to operate a particular software environment. A single agent is able to take actions in a variety of three-dimensional virtual worlds [6]. Recently, there have been a lot of studies that present the architecture of their agents and broader AI systems [7, 8]. LangChain analysed the cognitive architecture of agents [9]. However, these architectures often focus only on certain components or schemes. For instance, Packer et al. [7] explicitly covered the memory design of the agent. Andrew Ng briefly discussed reflection, tool use, planning and multi-agent collaboration [6]. Hassan et al. [10] demonstrated a high-level structure of FMware, consisting of Agentware and Promptware, and identified the challenges of software engineering for FMware. Gao et al. [11] proposed the roadmap for designing biomedical AI agents, and illustrated the required components.
We noticed that there is a lack of a holistic view of architecture design, making it challenging to develop FM-based agents. Moreover, software architecture comprises software elements, relations among them, and properties of both [12]. With the increase in the incorporation of machine learning into software and systems, methods to identify the impact on the reliability of machine learning are essential to ensure the reliability of the software and systems in which these algorithms reside [13]. Hence, frameworks that only list the high-level components supporting their functionality usually lack system-level thinking, with no explicit identification of software components, relationships among them, and their properties [3]. Our work presents 16 design patterns regarding agent goal-seeking and plan generation. The pattern catalogue can provide a holistic guidance to practitioners on the trade-off analysis for the design of foundation model-based agents.
@@ 3 Methodology
Based on the reported findings, we delved into the analysis of identified patterns for building integrating FM-based agents. Through the SLR, we have identified a series of architectural design challenges in the development and implementation of systems with the integration of agents. We then further conducted extensive review on this topic, which includes grey literature, and real-world applications (through scrutinising official websites and documents available), for identifying the known uses. Combining the findings of both SLR and the additional review, in this paper, we report our findings on 16 extracted patterns.
@@
Passive goal creator Analyse users’ articulated prompts through the dialogue interface to preserve interactivity, goal-seeking and intuitiveness.
@@
Prompt/response optimiser Optimise the prompts/responses according to the desired input or output content and format to provide standardisation, response accuracy, interoperability and adaptability.
@@
One-shot model querying Access the foundation model in a single instance to generate all necessary steps for the plan for cost efficiency, and simplicity.
Incremental model querying Access the foundation model at each step of the plan generation process to provide supplementary context, improve response accuracy and explainability
@@
Cross-reflection Use different agents or foundation models to provide feedback and refine the generated plan and reasoning process for better reasoning certainty, explainability, interoperability, inclusiveness, scalability and continuous improvement.
Human reflection Collect feedback from humans to refine the plan and reasoning process, to effectively align with human preference, improve contestability, effectiveness, fairness, and continuous improvement.
Voting-based cooperation Enable free opinions expression across agents and reach consensus by submitting their votes to improve diversity, efficient division of labor and fault tolerance.
Role-based cooperation Assign assorted roles, and finalise decisions in accordance with the roles of agents for better decision finalisation, division of labor, fault tolerance, scalability and accountability.
Debate-based cooperation Provide and receive feedback across multiple agents adjusts the thoughts and behaviors during the debate with other agents until a consensus is reached to improve decision finalisation, adaptability, explainability, response accuracy and critical thinking.
@@
Tool/agent registry Maintain a unified and convenient source to select diverse agents and tools to improve discoverability, efficiency and tool appropriateness.
@@ 4 Pattern Catalogue for Foundation Model-based Agents
A generated plan should be reviewed to ensure its accuracy, usability, completeness, etc. Self-refection, cross-reflection, and human reflection help the agent collect feedback from different reflective entities, and refine the plan and reasoning steps accordingly. Afterward, the agent can assign tasks to other traditional AI or non-AI systems, invoke external tools, and employ a set of agents for goal achievement by tool/agent registry. In particular, agents can work on the same task and finalise the results with voting-based, role-based, or debate-based cooperation, for instance, agents can act as different roles such as coordinator and worker. Multimodal guardrails can be applied to manage and control the inputs/outputs of foundation models. Please note that we omit the detailed architecture of agent-as-a-worker, and pattern application in several interactions for the simplicity of this diagram, for instance, the invocation of non-agent-AI/non-AI systems and external tools services can all apply tool/agent registry, and agent-as-a-worker can all have their respective datastore or knowledge base.
@@ 4.1 Passive Goal Creator
Problem: The agent may generate inaccurate responses to users’ instructions when given insufficient context.
@@
  • Interactivity. Users need to interact with the agent to provide instructions and receive responses.
  • Goal-seeking. The agent requires as much information as possible to understand users’ goals.
  • Intuitiveness. Simple tasks should be assigned and reported in an intuitive manner.
@@
Solution: Fig. 3 illustrates a simple graphical representation of passive goal creator. A foundation mode-based agent provides a dialogue interface where users can directly specify the context and problems, which are transferred to passive goal creator for goal generation. Meanwhile, the passive goal creator can also retrieve related information from memory, including the repository of artefacts being worked on, relevant tools used in recent tasks, conversation histories, and the positive and negative examples, which are appended to the user’s prompt for goal-seeking. The generated goals are sent to other components for further task decomposition and completion. In this case, the agent passively receives input from users and generates the strategies to achieve goals as it only receives the context information provided by users.
@@
  • Interactivity. Users can interact with the agent via a dialogue interface.
  • Goal-seeking. The agent can analyse user-provided context information to identify the objectives and create corresponding strategies.
  • Intuitiveness. Users can directly specify the goals, provide context to the agent, and receive the agent’s responses through the dialogue interface, which is intuitive and easy to use.
@@
  • Response inaccuracy. Users may have assorted backgrounds and experiences, while there are no standard prompt requirements. Unclear user-provided context information may result in hallucinated responses by the agent.
@@ 4.2 Proactive Goal Creator
  • Interactivity. Users are required to interact with the agent to provide instructions and receive responses.
  • Goal-seeking. The agent requires as much information as possible to understand users’ goals.
@@
Solution: Fig. 4 illustrates a simple graphical representation of proactive goal creator. In addition to the prompts received from dialogue interface, and relevant context retrieved from memory, the proactive goal creator can anticipate users’ goals by sending requirements to detectors, which will then capture and return the user’s surroundings for further analysis and comprehension to generate the goals, for instance, identifying the user’s gestures through cameras, recognising application UI layout via screenshots, etc. The proactive goal creator should notify users about context capturing and other relevant issues with a low false positive rate, to avoid unnecessary interruptions.
@@
  • Interactivity. Users can interact with the agent by providing multimodal information. In addition, in the multi-agent environment, proactive goal creator can facilitate the interactions between different agents by anticipating their decisions proactively.
  • Goal-seeking. The multimodal input can provide more detailed information for the agent to understand users’ goals, and increase the accuracy of generated responses.
@@ 4.3 Prompt/Response Optimiser
  • Response accuracy. Ensuring that prompts and responses are aligned with the ultimate goal or objective can facilitate the agent to achieve desired results.
@@
Solution: Fig. 5 illustrates a high-level graphical representation of prompt/response optimiser. A user may input initial prompts to the agent, however, such prompts may be ineffective due to the lack of relevant context, unintentional injection attacks, etc. In this regard, prompt/response optimiser can construct refined prompts and responses adhering to predefined constraints and specifications. These constraints and specifications outline the desired content and format for the inputs and outputs, ensuring alignment with the ultimate goal. A prompt or response template is often used in the prompt/response optimiser as a factory for creating specific instances of prompts or responses [20, 21]. This template offers a structured approach to standardise the queries and responses, improving the accuracy of the responses and facilitating interoperations with external tools or agents. For instance, a prompt template can contain the instructions to an agent, some examples for few-shot learning, and the question/goal for the agent to work.
@@
  • Response accuracy. The optimised prompts and responses adhere to user-defined conditions, hence they can achieve higher accuracy and relevance to the goals.
  • Interoperability. Interoperability between agent and external tools is facilitated by prompt/response optimiser, which can provide consistent and well-defined prompts and responses.
@@
  • Response inaccuracy. In certain cases, it may be difficult for prompt/response optimiser to capture and incorporate all relevant contextual information effectively, especially considering the ambiguity of users’ input, and dependency on context engineering. Consequently, the optimiser may struggle to generate appropriate prompts or responses.
@@ 4.4 Retrieval Augmented Generation (RAG)
  • Reasoning uncertainty. The reasoning process may be unreliable when the agent is required to accomplish domain-specific tasks that the agent has no such knowledge reserve.
@@
Solution: Fig. 6 illustrates a high-level graphical representation of retrieval augmented generation. RAG is a technique for enhancing the accuracy and reliability of agents with facts retrieved from other sources (internal or online data). The knowledge gaps where the agents are lacking in memory are filled using the parameterized knowledge generated in vector databases. For instance, during plan generation, specific steps may require information that is not within the original agent memory. The agent can retrieve information from the parameterized knowledge and use for planning, while the augmented response (i.e. plan) will be return back to the user. The retrieval process requires zero pretraining or fine-tuning of the model served by the agent which preserves the data privacy of local data, reduces training and computation costs, and also provides up-to-date and more precise information required. Currently there is a cluster of RAG techniques focusing on various enhancement aspects, data sources and applications, etc [22].
@@
  • Reasoning certainty. Agents can search and retrieve knowledge related to the given tasks, which ensure the reliability of reasoning steps.

### v2

@@ A Collection of Architectural Patterns
Foundation model-enabled generative artificial intelligence facilitates the development and implementation of agents, which can leverage distinguished reasoning and language processing capabilities to takes a proactive, autonomous role to pursue users’ goals. Nevertheless, there is a lack of systematic knowledge to guide practitioners in designing the agents considering challenges of goal-seeking (including generating instrumental goals and plans), such as hallucinations inherent in foundation models, explainability of reasoning process, complex accountability, etc. To address this issue, we have performed a systematic literature review to understand the state-of-the-art foundation model-based agents and the broader ecosystem. In this paper, we present a pattern catalogue consisting of 17 architectural patterns with analyses of the context, forces, and trade-offs as the outcomes from the previous literature review. The proposed catalogue can provide holistic guidance for the effective use of patterns, and support the architecture design of foundation model-based agents by facilitating goal-seeking and plan generation.
@@ 1 Introduction
  • Agents often struggle to fully comprehend and execute complex tasks, leading to the potential for inaccurate responses. This challenge may be intensified by the inherent reasoning uncertainties during plan generation and action procedures. For instance, across a long-term planning, the included steps may depend on each other, even slight deviation to a few steps can significantly impact the overall success rate.
  • Agents should not be entirely blamed for inaccurate response, since users may provide limited context, ambiguous goals or unclear instructions during the interaction with agents, which will result in underspecification [3, 4] in the reasoning process and response generation of agents.
  • The sophisticated internal architecture of agents and foundation models results in limited explainability, making them “black boxes” to stakeholders. Consequently, agents often struggle to interpret their reasoning steps, which can affect the reliability, robustness, and overall trustworthiness of agent systems.
  • The accountability process is complicated due to the interactions between various stakeholders, FM-based agents, non-agent AI models, and non-AI software applications within the overall ecosystem. Highly autonomous agents may delegate or even create other agents or tools for certain tasks. In this circumstance, responsibility and accountability may be intertwined among multiple entities.
In this regard, we present a catalogue of patterns for the design of foundation model-based agents in this paper. Please note that “agent” can be referred to i) AI acting on behalf of another entity, or; ii) AI that can take active roles or produces effect to achieve users’ goals. The former circumstance requires thorough analysis on governance perspective, while hereby, we claim that in this study, we focus on the second concept of “agents” that are capable of goal-seeking and plan generation. In software engineering, an architectural pattern is a reusable solution to a problem that occurs commonly within a given context in software design. Our pattern catalogue includes 17 patterns that were identified based on the study conducted by Lu et al. [5]. The intended audience of collected patterns is software architects and developers who are interested in FM-based agent design and implementation. The contributions of this paper include:
@@ 2 Background & Related Work
The introduction of OpenAI’s ChatGPT [6] in November 2022 has gathered over 100 million users in two months upon its release, becoming the fastest-growing consumer internet app of all time [7]. This has also initiated the race among big tech arms in the development of FM and GenAI products. For instance, Google rolled out its own GenAI product, the Bard models  [^3] https://ai.google/static/documents/google-about-bard.pdf, then released the Gemini  [^4] https://gemini.google.com/ as the updated version. Anthropic has also emerged as one of the major LLM providers since the launch of their Claude models  [^5] https://www.anthropic.com/claude There are also many open-sourced FMs, such as Llama  [^6] https://llama.meta.com/ and Mistral  [^7] https://mistral.ai/.
With the explosive growth of FMs, it is highly notable that FM-based agents come into the picture. AI agents are typically designed to operate a particular software environment. A single agent is able to take actions in a variety of three-dimensional virtual worlds [8]. Recently, there have been a lot of studies that present the architecture of their agents and broader AI systems [9, 10]. LangChain analysed the cognitive architecture of agents [11]. However, these architectures often focus only on certain components or schemes. For instance, Packer et al. [9] explicitly covered the memory design of the agent. Andrew Ng discussed reflection, tool use, planning and multi-agent collaboration [8] and provided a corresponding course [^8] https://www.deeplearning.ai/short-courses/ai-agentic-design-patterns-with-autogen/. Jain presented four design patterns for compound AI systems, including retrieval augmented generation, conversational AI, multi-agent communication, and co-pilot [12]. Yan et al. [13] shared their experience of building products with LLMs and the best practices they summarised during this process. Hassan et al. [14] demonstrated a high-level structure of FMware, consisting of Agentware and Promptware, and identified the challenges of software engineering for FMware. Gao et al. [15] proposed the roadmap for designing biomedical AI agents, and illustrated the required components.
We noticed that there is a lack of a holistic view of architecture design, making it challenging to develop FM-based agents. Moreover, software architecture comprises software elements, relations among them, and properties of both [16]. With the increase in the incorporation of machine learning into software and systems, methods to identify the impact on the reliability of machine learning are essential to ensure the reliability of the software and systems in which these algorithms reside [17]. Hence, frameworks that only list the high-level components supporting their functionality usually lack system-level thinking, with no explicit identification of software components, relationships among them, and their properties [5]. Our work presents 17 design patterns regarding agent goal-seeking and plan generation. The pattern catalogue can provide a holistic guidance to practitioners on the trade-off analysis for the design of foundation model-based agents.
@@ 3 Methodology
Based on the reported findings, we delved into the analysis of identified patterns for building integrating FM-based agents. Through the SLR, we have identified a series of architectural design challenges in the development and implementation of systems with the integration of agents. We then further conducted extensive review on this topic, which includes grey literature, and real-world applications (through scrutinising official websites and documents available), for identifying the known uses. Combining the findings of both SLR and the additional review, in this paper, we report our findings on 17 extracted patterns.
@@
Passive goal creator Analyse users’ articulated prompts through the dialogue interface to preserve interactivity, goal-seeking and efficiency.
@@
Prompt/response optimiser Optimise the prompts/responses according to the desired input or output content and format to provide standardisation, goal alignment, interoperability and adaptability.
@@
One-shot model querying Access the foundation model in a single instance to generate all necessary steps for the plan for cost efficiency and simplicity.
Incremental model querying Access the foundation model at each step of the plan generation process to provide supplementary context, improve reasoning certainty and explainability
@@
Cross-reflection Use different agents or foundation models to provide feedback and refine the generated plan and reasoning process for better reasoning certainty, explainability, inclusiveness and scalability.
Human reflection Collect feedback from humans to refine the plan and reasoning process, to effectively align with human preference, improving contestability and effectiveness.
Voting-based cooperation Enable free opinions expression across agents and reach consensus by submitting their votes to preserve fairness, accountability and collective intelligence.
Role-based cooperation Assign assorted roles, and finalise decisions in accordance with the roles of agents for to facilitate division of labor, fault tolerance, scalability and accountability.
Debate-based cooperation Provide and receive feedback across multiple agents adjusts the thoughts and behaviors during the debate with other agents until a consensus is reached to improve adaptability, explainability and critical thinking.
@@
Tool/agent registry Maintain a unified and convenient source to select diverse agents and tools to improve discoverability, efficiency, tool appropriateness and scalability.
Agent evaluator Perform testing to assess the agent regarding diverse requirements and metrics, ensuring the functional suitability, adaptability with improved flexibility.
@@ 4 Pattern Catalogue for Foundation Model-based Agents
A generated plan should be reviewed to ensure its accuracy, usability, completeness, etc. Self-refection, cross-reflection, and human reflection can help the agent to collect feedback from different reflective entities, and refine the plan and reasoning steps accordingly. Afterwards, the agent can assign tasks to other narrow AI-based or non-AI systems, invoke external tools, and employ a set of agents for goal achievement by tool/agent registry. In particular, agents can work on the same task and finalise the results with voting-based, role-based, or debate-based cooperation. For instance, agents can act as different roles such as coordinator and worker. Multimodal guardrails can be applied to manage and control the inputs/outputs of foundation models. Please note that we omit the detailed architecture of agent-as-a-worker, and pattern application in several interactions for the simplicity of this diagram, for instance, the invocation of non-agent-AI/non-AI systems and external tools services can all apply tool/agent registry, and agent-as-a-worker can all have their respective datastore or knowledge base. Finally, we claim that developers can evaluate the performance of agents at both design-time and runtime via agent evaluator.
@@ 4.1 Passive Goal Creator
Context: When querying agents to address certain issues, users provide related context and explain the goals in prompts.
Problem: Users may lack expertise of interacting with agents, and the provided information can be ambiguous for goal achievement.
@@
  • Underspecification. Users may not be able to provide thorough context information and specify precise goals to agents.
  • Efficiency. Users expect quick responses from agents.
@@
Solution: Fig. 3 illustrates a simple graphical representation of passive goal creator. A foundation model-based agent provides a dialogue interface where users can directly specify the context and problems, which are transferred to passive goal creator for goal determination. Meanwhile, the passive goal creator can also retrieve related information from memory, including the repository of artefacts being worked on, relevant tools used in recent tasks, conversation histories, and the positive and negative examples, which are appended to the user’s prompt for goal-seeking. The generated goals are sent to other components for further task decomposition and completion. In this case, the agent passively receives input from users and generates the strategies to refine and clarify users’ goals, as it only receives the context information directly provided by users. Please note that in multi-agent systems, an agent can send prompts by invoking the API of another agent to assign specific task, while the latter agent analyses the received information and determine the goal.
@@
  • Interactivity. Users or other agents can interact with an agent via a dialogue interface or related APIs.
  • Goal-seeking. The agent can analyse user-provided context and retrieve related information from memory, to identify and determine the objectives and create corresponding strategies.
  • Efficiency. Users can directly send prompts to the agent through the dialogue interface, which is intuitive and easy to use.
@@
  • Reasoning uncertainty. Users may have assorted backgrounds and experiences. Unclear or ambiguous context information may intensify the reasoning uncertainties, especially considering there are no standard prompt requirements.
@@ 4.2 Proactive Goal Creator
  • Underspecification. i) Users may not be able to provide thorough context information and specify precise goals to agents. ii) Agents can only retrieve limited information from the memory.
@@
Solution: Fig. 4 illustrates a simple graphical representation of proactive goal creator. In addition to the prompts received from dialogue interface, and relevant context retrieved from memory, the proactive goal creator can anticipate users’ goals by sending requirements to detectors, which will then capture and return the user’s surroundings for further analysis and comprehension to generate the goals, for instance, identifying the user’s gestures through cameras, recognising application UI layout via screenshots, etc. The proactive goal creator should notify users about context capturing and other relevant issues with a low false positive rate, to avoid unnecessary interruptions. In addition, the captured environment information can be stored in the agent’s memory (or knowledge base) to establish “world models” [21, 22] to improve its ability to comprehend the real world.
@@
  • Interactivity. An agent can interact with users or other agents by anticipating their decisions proactively with captured multimodal context information.
  • Goal-seeking. The multimodal input can provide more detailed information for the agent to understand users’ goals, and increase the accuracy and completeness of goal achievement.
@@ 4.3 Prompt/Response Optimiser
  • Goal alignment. Ensuring that prompts and responses are aligned with the ultimate goal or objective can facilitate the agent to achieve desired results.
@@
Solution: Fig. 5 illustrates a high-level graphical representation of prompt/response optimiser. A user may input initial prompts to the agent, however, such prompts may be ineffective due to the lack of relevant context, unintentional injection attacks, redundancy, etc. In this regard, prompt/response optimiser can construct refined prompts and responses adhering to predefined constraints and specifications. These constraints and specifications outline the desired content and format for the inputs and outputs, ensuring alignment with the ultimate goal. A prompt or response template is often used in the prompt/response optimiser as a factory for creating specific instances of prompts or responses [26, 27]. This template offers a structured approach to standardise the queries and responses, improving the accuracy of the responses and facilitate their interoperations with external tools or agents. For instance, a prompt template can contain the instructions to an agent, some examples for few-shot learning, and the question/goal for the agent to work.
@@
  • Goal alignment. The optimised prompts and responses adhere to user-defined conditions, hence they can achieve higher accuracy and relevance to the goals.
  • Interoperability. Interoperability between agent and external tools is facilitated by prompt/response optimiser, which can provide consistent and well-defined prompts and responses for task execution.
@@
  • Underspecification. In certain cases, it may be difficult for prompt/response optimiser to capture and incorporate all relevant contextual information effectively, especially considering the ambiguity of users’ input, and dependency on context engineering. Consequently, the optimiser may struggle to generate appropriate prompts or responses.
@@ 4.4 Retrieval Augmented Generation (RAG)
  • Lack of knowledge. The reasoning process may be unreliable when the agent is required to accomplish domain-specific tasks that the agent has no such knowledge reserve.
@@
Solution: Fig. 6 illustrates a high-level graphical representation of retrieval augmented generation. RAG is a technique for enhancing the accuracy and reliability of agents with facts retrieved from other sources (internal or online data). The knowledge gaps where the agents are lacking in memory are filled using the parameterized knowledge generated in vector databases. For instance, during plan generation, specific steps may require information that is not within the original agent memory. The agent can retrieve information from the parameterized knowledge and use for planning, while the augmented response (i.e. plan) will be return back to the user. The retrieval process requires zero pretraining or fine-tuning of the model served by the agent which preserves the data privacy of local data, reduces training and computation costs, and also provides up-to-date and more precise information required. The retrieved local data can be sent back to the agent via prompts (need to consider the context window size), whereas the agent is able to process the information and generate plans via in-context learning. Currently there is a cluster of RAG techniques focusing on various enhancement aspects, data sources and applications [28], for instance, federated RAG [29], graph RAG [30], etc.
@@
  • Knowledge retrieval. Agents can search and retrieve knowledge related to the given tasks, which ensures the reliability of reasoning steps.